I need help solving systems
-3x + 6y = 24
2x + y =-1


and

-3x + y =5
2x - y =5

Answer by Maths68(1474)

-3x + 6y = 24...........(1)
2x + y =-1..............(2)
Multiply (1) by 2 and (2) by 3 and add
-6x+12y=48
6x+3y=-3
-------------------
15y=45
15y/15=45/15
y=3
Put the value of y in (1)
-3x + 6y = 24...........(1)
-3x + 6(3) = 24
-3x + 18 = 24
-3x = 24-18
-3x = 6
-3x/-3 = 6/-3
x=-2
x,y=-2,3


Check
======
Put the value of x and y in (1)
-3x + 6y = 24...........(1)
-3(-2) + 6(3) = 24
6+18=24
24=24
Put the value of x and y in (2)
2x + y =-1..............(2)
2(-2)+3=-1
-4+3=-1
-1=-1
============================================================================================

-3x + y =5...............(1)
2x - y =5................(2)
Add (1) and (2)
-3x + y =5
2x - y =5
-----------------
-x=10
x=-10
Put the value of x in (1)
-3x + y =5...............(1)
-3(-10) + y =5
30+y=5
y=5-30
y=-25
x,y=-10,-25


Check
======
Put the values of x and y in (1)
-3x + y =5...............(1)
-3(-10)+(-25)=5
30-25=25
5=5
Put the values of x and y in (2)
2x - y =5................(2)
2(-10) - (-25) =5
-20+25=5
5=5